OpenStudy (anonymous):

determine the equation of the axis symmentry. -5x^2-x+9

OpenStudy (anonymous):

\[x=-\frac{b}{2a}\] which in your case is \(x=-\frac{-1}{2\times (-5)}=-\frac{1}{10}\)

OpenStudy (anonymous):

thank you (: